Opened 12 years ago

Last modified 11 years ago

#464 new flight processing

GB12/05, flight day 259/2012, Monks Wood

Reported by: knpa Owned by:
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by edfi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/GB12_05-2012_259_Monks_Wood

Data arrived from ARSF via SATA disk on 24/09/2012

GB12/05 Scientific objective: unknown

GB12/05 Priority: a4m

PI: Ross Hill

Sensors:

Camera (13/12/12)
Eagle (09/04/13)
Hawk (09/04/13)
LiDAR (16/01/13)
FW LiDAR (09/04/13)

Change History (41)

comment:1 Changed 12 years ago by knpa

  • Summary changed from GB12/05, flight day 249/2012, Monks Wood to GB12/05, flight day 259/2012, Monks Wood

comment:2 Changed 12 years ago by besm

Navigation processed.

comment:3 Changed 12 years ago by besm

Eagle and hawk raw files are listed as SWIR259a-*, but the flight has been renamed from 259a to 259. Hyperspectral processing doesn't work. I'm renaming them to SWIR259-* now.

comment:4 Changed 12 years ago by anhi

RCD Deliver Created

Version 0, edited 12 years ago by anhi (next)

comment:5 Changed 12 years ago by edfi

Beginning Lidar

comment:6 Changed 12 years ago by anhi

  • Description modified (diff)

comment:7 Changed 12 years ago by anhi

Delivered Camera to Ross Hill sent on 13th December.

comment:8 Changed 12 years ago by besm

Hyperspectral delivery proceeding without vectors. Delivery created and readme is ongoing.

comment:9 Changed 12 years ago by besm

Readme created. Ready to delivery check.

Please make sure to check all files are in the right place and appropriately named, as the make delivery script may have had a silent problem.

comment:10 Changed 12 years ago by edfi

Started D/C of Hyperspectral

comment:11 Changed 12 years ago by edfi

Only three bands processed needs re-processing

comment:12 Changed 12 years ago by besm

Starting reprocessing.

comment:13 Changed 12 years ago by besm

Finished reprocessing. Ready to re-delivery check.

comment:14 Changed 11 years ago by edfi

Started Hyperspectral delivery check

comment:15 Changed 11 years ago by edfi

Hyperspectral check complete was missing log sheet which has now been added. Everything is fine. Currently zipping mapped files

comment:16 Changed 11 years ago by edfi

  • Description modified (diff)

comment:17 Changed 11 years ago by edfi

LiDAR delivery created

comment:18 Changed 11 years ago by anhi

LiDAR Delivery Check:

  • Hand classify the electricity pylons/wires down the left side of the flight, around x:515758.

No other problems

comment:19 Changed 11 years ago by edfi

Power lines have been manually reclassified,re-running delivery

comment:20 Changed 11 years ago by anhi

LiDAR Delivery Check:

No problems.

comment:21 Changed 11 years ago by edfi

Vector email sent to PI

comment:22 Changed 11 years ago by edfi

LiDAR sent 16/01/13

comment:23 Changed 11 years ago by edfi

  • Description modified (diff)

comment:24 Changed 11 years ago by edfi

Started FW processing

comment:25 Changed 11 years ago by edfi

Starting OS vector check

comment:26 Changed 11 years ago by edfi

Vector screen-shots added and Readme updated for LiDAR and Hyperspectral ready for update check

comment:27 Changed 11 years ago by anhi

Lidar vector check:

Screenshot looks good but no Readme in delivery.

comment:28 Changed 11 years ago by edfi

Readme moved into right delivery folder

comment:29 Changed 11 years ago by anhi

LiDAR Readme check:

No problems found.

comment:30 Changed 11 years ago by anhi

Hyperspectral vector check:

  • Readme aplmap command bandlist should be ALL

Screenshots and rest of readme look fine.

comment:31 Changed 11 years ago by edfi

aplmap command updated to all. Ready for check

comment:32 Changed 11 years ago by anhi

Hyperspectral vector check:

  • gdal_translate example command in Readme needs correcting

comment:33 Changed 11 years ago by edfi

Hyperspectral ReadMe remade with corrected GDAL commands

comment:34 Changed 11 years ago by anhi

Hyperspectral readme check:

No problems

comment:35 Changed 11 years ago by edfi

FW delivery created by anhi, checked by edfi

comment:36 Changed 11 years ago by edfi

Reprocessing Eagle with December 2012 calibration

comment:37 Changed 11 years ago by edfi

Eagle reprocessed and new delivery and ReadMe made. Ready for delivery check

comment:38 Changed 11 years ago by tipo

Hyperspectral delivery checked, eagle viewvector files were missing, fixed.

comment:39 Changed 11 years ago by edfi

  • Description modified (diff)

comment:40 Changed 11 years ago by edfi

Hyperspectral (December 2012 Calibration) and FW LiDAR dispatched 09/04/13 to PI

comment:41 Changed 11 years ago by tipo

Prepared for archiving (25/9/2013)

Note: See TracTickets for help on using tickets.